What is the potential difference required to stop an electron traveling at an initial speed of 2.29 x 107 m/s? (mass of an electron is 9.1 X 10-31 kg)
1
Expert's answer
2020-02-28T10:20:48-0500

E = mv2 / 2 = 9.1 × 10-31 × (2.29 × 107)2 / 2 = 2.39 × 10-16 (J) = 2.39 × 10-16 / 1,602 × 10−19 = 14.3 eV.

Answer: 14.3 eV.